OpenStudy (anonymous):

if the area of circle is halved and the radius is decreased by n then the radius will be

OpenStudy (ajprincess):

Let the initial radius of the circle be r and area be \(\pi r^2\) let the new radius be \(r_1\) \(r_1=r-n\) new area is \(\Large\frac{\pi r^2}{2}\) \(\Large\frac{\pi r^2}{2}=\pi {r_1}^2\) \(r^2=2{r_1}^2\) \(r^2=2(r-n)^2\) solve for r.Does that help?

OpenStudy (anonymous):

I wonder if n was a factor (such as 0.7) rather than a number to be subtracted.

OpenStudy (anonymous):

Making the question r2 = n r1 when A2 = (1/2) A1

OpenStudy (anonymous):

Glad to have helped.

OpenStudy (anonymous):

thanks I need the answer as n(2+root2)

OpenStudy (ajprincess):

if so then u need to solve the equation I gave. \(r^2=2(r^2-2rn+n^2)\) \(r^2=2r^2-4rn+2n^2\) \(r^2-4rn+2n^2=0\) use \[r=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] here b=-4n c=2n^2 a=1